OpenStudy (anonymous):

lisa uses 3 pounds of sugar to make 100 cookies. How many cookies can she make with 5 pounds of sugar. round to the nearest whole cookie

OpenStudy (jdoe0001):

\(\large \begin{array}{llll} sugar&cookies\\ 3&100\\ 5&x \end{array}\implies \cfrac{3}{5}=\cfrac{100}{x}\)

OpenStudy (texaschic101):

set up a proportion 3/100 = 5/x (3 lbs sugar to 100 cookies = 5 lbs to x cookies) cross multiply (3)(x) = (100)(5) 3x = 500 x = 500/3 x = 166.66 = 167 cookies